Find the volume of the region bounded above by the plane z = y/2 and below by the rectangle.

R ∶ 0 ≤ x ≤ 4,0 ≤ y ≤ 2


1
Expert's answer
2022-03-20T06:41:39-0400

V= "\int_0^4 dx \int_0^2 dy \int_0^{y/2}dz= \int_0^4 dx \int_0^2 \frac y2 dy ="

"\int_0^4 dx \frac{y^2}{4}|_0^2= \int_0^4 dx =4"

Answer: V=4.